OpenStudy (anonymous):

if angle ab is complementary w/ angle bc and ab is equal to 56 degrees what is the measure of angle bc?

OpenStudy (anonymous):

X+8X=90 9X=90 X=90/9 X=10 DEGREES FOR THE SMALLER ANGLE. 8*10=80 DEGREES FOR THE LARGER ANGLE. PROOF: 10+80=90 90=90

OpenStudy (anonymous):

can draw a tringle

OpenStudy (anonymous):



OpenStudy (anonymous):

what is this

OpenStudy (anonymous):

what is what???